OpenStudy (anonymous):

a roller coaster at six flags is 1.5 kilometers long. how many miles is it?

OpenStudy (phi):

do you know how many miles in a kilometer or km in a mile?

OpenStudy (anonymous):

1 mile = 1.6 km

OpenStudy (phi):

to do these kinds of problems you should first find a ratio of miles to km for example \[ \frac{1 \text{ mile}}{1.6 \text{ km}}\] then you multiply so that the units cancel \[ 1.5 \cancel{\text{ km}}\cdot \frac{1 \text{ mile}}{1.6 \cancel{\text{ km}}}\]

OpenStudy (phi):

google can 1.609344 kilometers in 1 mile most people round this to 1.6 km per mile once you know this ratio, look at the problem you are working on. you want to convert km to miles that means you want to multiply \[ km \cdot \frac{mile}{km}\] so that the km cancel, and you are left with mile 1.6 km per mile means the same as 1.6 km / 1 mile or \(\frac{1.6 \text{ km}}{1 \text{ mile}}\) notice that this ratio is upside down from what you want, so flip it to get the correct ratio \[ \frac{1 \text{ mile}}{1.6 \text{ km}}\] now use it to do your conversion (see above)

OpenStudy (phi):

you get \[ \frac{1.5}{1.6} \text{ miles} \] which you can simplify